OpenStudy (anonymous):

doing trigs.. prove that \frac {<\[\sin A\]>}{<\[1 + \cos A\]>} + \frac {<\[1 + \cos A\]>}{<\[\sin A\]>} = \[2\csc A\] oh my god I hope that turns out right. derp.

OpenStudy (anonymous):

......... IT DID NOOOOOOOT

OpenStudy (dumbcow):

lol

OpenStudy (dumbcow):

\[\frac{\sin A}{1+\cos A}+\frac{1+\cos A}{\sin A} =2\csc A\]

OpenStudy (dumbcow):

is that it?

OpenStudy (dumbcow):

anyway if you combine the fractions on left side and use trig identity sin^2 + cos^2 = 1 then it should work out nicely